OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

15x-5y = 30 -5y = 30-15x y = 3x-6 y = 3(0)-6 y = 0-6 y = -6, (0,6) (0) = 3x-6 0 = 3x-6 -3x = -6 x = 2, (2,0) Solution: x-intercept = (2,0); y-intercept = (0,-6)